• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Camsonne

    Log entry time 22:56:58 on November 8,2006

    Entry number 189103

    keyword=Lower current running 55 uA for Compton Measurement

    I asked to reduce our current with the slits to reduce the background for best signal over background, background ate now down to about 80 Hz as expected for about 250 Hz of rate with laser on

    FIGURE 1